OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, find h[j(2)]. 11 −11 13 −13

OpenStudy (akashdeepdeb):

j(2) = -2(2) = -4 right? And h(x) = 3x-1 So h(j(2)) = h(-4) = 3(-4) - 1 = -12-1 = -13 Understood? :)